The following arbitrary measurements are made and the errors sited are the maximum errors A=15.21±0.01,B=10.82±0.05, C=11.00±0.03 . If D=A+B+C; 
a) calculate the maximum error in D
b) if the errors sited are standard errors,calculate the standard error in D. 


1
Expert's answer
2021-08-16T08:36:32-0400

a)


"\Delta D=0.01+0.05+0.03=0.09"

b)


"SE_D=\sqrt{0.01^2+0.05^2+0.03^2}=0.06"